OpenStudy (anonymous):

Tan x=1.8040. Find x.

OpenStudy (anonymous):

Use the inverse tangent to get 61 degrees (I assume you need the angle measure) so tan^-1=60.99935688

OpenStudy (anonymous):

sorry I meant to type tan^-1 (1.8040)=60.99935688

OpenStudy (shubhamsrg):

well we have sin x = 1.8 cos x sq. both sides sin^2 x = 3.24 cos^2 x 1 - cos^2 x = 3.24 cos^2 x cos^2 x = 1/4.24 => cos x = 1/2.05, -1/2.05 (approx) thus x is 2n(pi)+ pi/6 ,, approximately

myininaya (myininaya):

f(x)=tan(x) is an odd function f(-x)=-f(x) -f(-x)=f(x) So this means -tan(-x)=tan(x) So we have two equations tan(x)=1.804 -tan(-x)=1.804 Let n be an integer But we also have that tan(x)=tan(x+n*pi) So we also have -tan(-x)=-tan(-x+n*pi) So we can rewrite those two equations as tan(x+n*pi)=1.804 -tan(-x+n*pi)=1.804 Now this means when we have tan(x+n*pi)=1.804 tan(-x+n*pi)=-1.804 Now arctan( ) of both sides and blah blah